A rod, 1 cm2 x 200 cm and mass 2 kg, is clamped at its center. When vibrating
longitudinally it emits its third overtone in unison with a tuning fork making 6000 vps.
How much will the rod be elongated if, when clamped at one end, a stretching
force of 980 N is applied to the other end?
Answer #1

mass volume 2 a Z tes A XL v=1cm² x 200cm 200 cm3 2x 10-4 m3 200 cm3 - os 1x 10 4 m3 from eq? = r2 x a Y = (3428.40)? x 104 T980 X 20 X 10-2 AL = 1.175 X lol1 x ALT 980 x2 1.175 X 16+7 0.17 mm OR AL = 1.7 x 10-4